OpenStudy (anonymous):

The angle between 0 and 2π in radians that is coterminal with the angle 34π/5 in radians is

OpenStudy (anonymous):

Since π radians is 180 degrees, multiply the number of π radians by 180: (34/5)(180) = 1224 degrees In order to find a coterminal angle between 0 and 2π (between 0 and 360 degrees), continually subtract 360 degrees from this answer until you reach a value less than 360: 1224 - 360 = 864 - 360 = 504 - 360 = 144 Now that the coterminal angle between 0 and 360 degrees has been found, convert back to radians by dividing this angle by 180, and multiplying the result by π radians: (144/180) = (4/5) x π = 4π/5